Celia bought 12 apples. She ate two of them. She sold the remaining apples for 20 cents more than she bought each of them. She profited $1.00. How much did she sell the apples for?
Answer by checkley71(8403)

10(X+.20)=12X+1 OR 10X+2=12X+1 OR 2X=1 OR X=.50 PROOF
10(.50+.20)=12*.5+1 OR 10(.70)=6+1 OR 7=7 SHE SOLD THE APPLES FOR 70 CENTS.